### v1

# Novel View Synthesis using DDIM Inversion

Abstract
Synthesizing novel views from a single input image is a challenging task. It requires extrapolating the 3D structure of a scene while inferring details in occluded regions, and maintaining geometric consistency across viewpoints. Many existing methods must fine‑tune large diffusion backbones using multiple views or train a diffusion model from scratch, which is extremely expensive. Additionally, they suffer from blurry reconstruction and poor generalization. This gap presents the opportunity to explore an explicit lightweight view translation framework that can directly utilize the high-fidelity generative capabilities of a pretrained diffusion model while reconstructing a scene from a novel view. Given the DDIM-inverted latent of a single input image, we employ a camera pose-conditioned translation U-Net, TUNet, to predict the inverted latent corresponding to the desired target view. However, the image sampled using the predicted latent may result in a blurry reconstruction. To this end, we propose a novel fusion strategy that exploits the inherent noise correlation structure observed in DDIM inversion. The proposed fusion strategy helps preserve the texture and fine-grained details. To synthesize the novel view, we use the fused latent as the initial condition for DDIM sampling, leveraging the generative prior of the pretrained diffusion model. Extensive experiments on MVImgNet demonstrate that our method outperforms existing methods. Code is available.
(a) MvImgNet test set
(b) Out-of-domain real-world images
Figure 1: (a) High-resolution ($512\times 512$) novel-view synthesis on the MvImgNet test set from a single input image and camera parameters, (b) Zero-shot synthesis on out-of-domain images downloaded from Unsplash.
Introduction
Novel view synthesis is a fundamental task in computer vision and graphics. Remarkable works such as NeRFs (Mildenhall et al. 2021) and 3DGS (Kerbl et al. 2023) are extensively used in 3d scene understanding. Several works improve upon these foundational works. However, their dependence on scene-level optimization and the need for a dense set of views limit usability. Diffusion models (Rombach et al. 2022; Podell et al. 2023), have gained significant traction for the task of novel view synthesis (Elata et al. 2025; Tang et al. 2025). A classical approach is to fine-tune these models on 3D datasets along with a module that encodes the 3D geometry into the architecture (Tang et al. 2023; Liu et al. 2024; Long et al. 2024; Huang et al. 2024; Bourigault and Bourigault 2024; Gao et al. 2024). However, the generated outputs lack consistency in multiview reconstruction as the generation is not entirely controllable and results in images of inadequate quality, and often creates blurry results for long-range viewpoint reconstruction.
DDIM (Song, Meng, and Ermon 2020) proposed a deterministic inversion “DDIM Inversion”, which sequentially adds noise to an image to obtain a noisy latent. The noisy latent can be retraced to the original image using DDIM sampling. This latent encapsulates the signal and the noise that contribute to the mean and variance, which changes the distribution of the noise latent at each inversion time. Previous works such as (Garibi et al. 2024; Mokady et al. 2023) try to optimize or configure this noise representation to better align with the given task. (Staniszewski, Kuciński, and Deja 2025) study inversion noise in detail and claim that the DDIM inversion latent space is less manipulative, which makes direct interpolation with this noise latent difficult for tasks such as novel view synthesis and editing.
This paper proposes a method to generate a novel view from a given input image and camera parameters. Our pipeline works entirely in the DDIM-inverted latent space. We first learn to map an input view latent to a target latent using a translation U‑Net called TUNet. This mapping only approximates a coarse-grained version of the target view. This is due to the fact that diffusion models exhibit spectral bias and favor low-frequency components (Choi et al. 2022). In order to induce the high frequency components, we introduce a novel noisy latent fusion strategy. Notably, we use pretrained diffusion model, and only train a lightweight latent‑space translation network, TUNet, for view transformation. We perform extensive experiments in diverse settings, and show that our work extends to unseen categories as well as out of domain images obtained from the web. Sample results are shown in Figure 1. We claim the following key contributions:
  • We propose a method for translation of input DDIM-inverted latent to a target latent. The target latent can be decoded by a pretrained diffusion model’s VAE decoder to obtain the target novel view.
  • The translated latent may only result in a coarse-grained image with the broad structure of the target image being preserved. In order to inject high frequency details, we propose a novel fusion strategy. TUNet’s coarse output is fused with the high-variance noise obtained from our fusion strategy. The fused latent can be used to initialize DDIM sampling, which reconstructs a high‑quality novel view with consistent geometry and vivid fine‑grained detail.
  • In our experiments, we show that the method achieves superior results in terms of LPIPS, PSNR, SSIM, and FID.
Related Work
Neural Radiance Field: Neural field approaches, such as Neural Radiance Fields (NeRF) (Mildenhall et al. 2021), use learnable functions to map 3D spatial coordinates and viewing directions to volumetric density and color. These models synthesize novel views by performing volumetric rendering via ray marching through the learned scene representation. NeRF has demonstrated that high-quality novel views can be rendered when trained on a dense set of input views.
While recent extensions such as PixelNeRF (Yu et al. 2021), IBRNet (Wang et al. 2021), MultiDiff (Müller et al. 2024), and others (Henzler et al. 2021; Liu et al. 2022; Wu et al. 2023) aim to perform view synthesis from fewer input views, they often suffer in regions with missing or occluded content. Because these models make deterministic predictions without explicit uncertainty modeling, the generated output tends to average over ambiguities, leading to blurry and less plausible reconstruction in unobserved regions.
Gaussian Splatting: 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023; Peng et al. 2024; Zheng et al. 2025) represent scenes using a set of anisotropic 3D Gaussians. Gaussian Splatting methods are deterministic and depend heavily on accurate multi-view geometry or densely sampled camera poses (Li et al. 2024). When applied in sparse-view or single-view settings, they often fail to generate plausible content in unseen regions because they lack generative priors.
In contrast, our work targets novel view synthesis given only a single input image and a target camera pose. This setting spans both short and long-range viewpoint changes. Under such conditions, methods such as NeRF (Mildenhall et al. 2021) and 3DGS (Kerbl et al. 2023) struggle to extrapolate effectively from a single image, even when augmented with generative guidance as in (Szymanowicz et al. 2024; Tewari et al. 2023).
Transformers: NViST (Jang and Agapito 2024) adopts a transformer-based encoder-decoder architecture (Vaswani et al. 2017; Dosovitskiy et al. 2021) to predict a radiance field from a single image, enabling novel view synthesis via NeRF-style volumetric rendering. However, NViST suffers from loss of fine details due to aggressive downsampling (by a factor of 12), and it struggles to synthesize long-range viewpoints (when the target frame is more than 15 frames away from the input frame in a 30‑frame sweep).
Diffusion models: Diffusion models can be leveraged to generate plausible content in the unobserved regions of the input views. In the following, we identify them as the ones which finetune pre-trained diffusion models, or train diffusion models from scratch.
Pretrained Diffusion Models: MVDiffusion (Tang et al. 2023), Zero123++ (Shi et al. 2023), SyncDreamer (Liu et al. 2024), Wonder3D (Long et al. 2024), EpiDiff (Huang et al. 2024), BoostDream (Yu et al. 2024), MVDiff (Bourigault and Bourigault 2024), CAT3D (Gao et al. 2024), Magic-Boost (Yang et al. 2024), Cycle3D (Tang et al. 2025), use a pretrained or finetuned diffusion model. MVDiffusion modifies the Stable Diffusion architecture by introducing a cross-branch attention mechanism, known as correspondence-aware attention (CAA), to model inter-view dependencies. SyncDreamer constructs a view frustum feature volume from all the target noisy views and injects these into the pretrained denoising Unet using depth-wise attention layers. EpiDiff (Huang et al. 2024) integrates an attention module guided by epipolar constraints into the intermediate and decoding stages of the U-Net, enabling the model to capture generalized epipolar geometry across views.
Most existing approaches either fine-tune the diffusion model or inject spatial features corresponding to the target view into the base model’s denoising U-Net. Such features are typically derived from volumetric projections or depth estimates. However, models that follow this paradigm often struggle with scene-level reconstructions and are usually trained on object-specific datasets, which may limit the generalization to complex scenes.
In contrast, our method does not modify or inject any learned features into the U-Net of the diffusion model. Instead, we provide external conditioning input to TUNet to obtain the latent corresponding to the target view.
The latent text space of diffusion models has also been explored for view synthesis Burgess, Wang, and Yeung-Levy (2024). However, it is typically limited to small-scale datasets and is not directly applicable to scene-level novel view synthesis tasks.
Figure 2: Overview: Given a single reference image $\mathbf{x_{\text{ref}}}$, we first apply DDIM inversion up to $t=600$ to obtain the mean latent $\mathbf{z}_{\text{ref},\mu}^{\text{inv}}$. This, together with camera intrinsics/extrinsics, class embeddings, and ray information, is fed into our translation network TUNet. TUNet predicts the target-view mean latent $\tilde{\mathbf{z}}_{\text{tar},\mu}^{\text{inv}}$, which we combine with the corresponding noise component via one of our fusion strategies to form the initial DDIM latent $\tilde{\mathbf{z}}_{tar}^{\text{inv}}$. Finally, this latent is sampled by a pre-trained diffusion model to synthesize the novel view image.
Training Diffusion Model from Scratch: Several recent works train diffusion models from scratch for novel view synthesis, including Tseng et al. (2023), Photometric-NVS (Yu et al. 2023a), DiffDreamer (Cai et al. 2023), GIBR (Anciukevičius et al. 2024), and (Henderson et al. 2024). Photometric-NVS (Yu et al. 2023a) introduces a two-stream latent diffusion architecture that independently processes the source and noisy target views, while exchanging information via pose-conditioned cross-attention mechanisms. GIBR (Anciukevičius et al. 2024) models 3D scenes using IB-planes and trains the diffusion process directly in pixel space, enabling learning of a joint distribution over multi-view observations and camera poses.
Training entire diffusion models end-to-end is computationally expensive and requires large-scale datasets to achieve high-resolution and photorealistic reconstruction. In contrast, our method operates in the DDIM-inverted latent space at a fixed timestep, which corresponds to a weak yet informative signal. This allows us to perform translation from a given latent to a target latent using a lightweight translation U-Net. Operating in the latent space significantly simplifies the view translation task, as the model works with compact, semantically rich representations rather than raw pixels. Our fusion strategy provides the necessary information regarding the high-frequency scene details. The final novel view is synthesized using a pretrained diffusion pipeline, which decodes the predicted latent.
Method
Given a single reference image and camera parameters of the target viewpoint, our work addresses the task of novel view synthesis. Inspired by the deterministic behavior of DDIM inversion, we perform view synthesis entirely in the DDIM-inverted latent space. A dedicated translation network, TUnet, is trained to map the source latent to the target latent corresponding to the novel viewpoint. To induce the high frequency scene details, we propose a fusion strategy. The resulting latent is then passed through a pretrained diffusion model to generate the final high-fidelity novel view. Our method is illustrated in Figure 2.
Spectral Behavior of Diffusion
In (Kingma and Gao 2023; Falck et al. 2025), authors study the spectral behavior of diffusion. The forward diffusion (Ho, Jain, and Abbeel 2020) process is given by:
$\mathbf{x}_{t}=\underbrace{\sqrt{\overline{\alpha}_{t}}\,\mathbf{x}_{0}}_{\text{signal}}+\underbrace{\sqrt{1-\overline{\alpha}_{t}}\,\boldsymbol{\epsilon}}_{\text{noise}},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $\mathbf{x}_{0}$ is the clean latent, $\mathbf{x}_{t}$ is the latent corresponding to the timestep $t$ and $\bar{\alpha}_{t}$ is the scaling factor. High-frequency components, representing fine details, are degraded more rapidly and prior to low-frequency components during the forward diffusion process (Kingma and Gao 2023; Falck et al. 2025). This property is also consistent in the reverse process of diffusion. As shown in Choi et al. (2022), diffusion models inherently favor lower frequencies, which implies that more emphasis must be placed on modelling high-frequency details. In addition, the noise component is often observed to deviate from a standard multivariate Gaussian distribution (Staniszewski, Kuciński, and Deja 2025). At later iterations of inversion, the noise encapsulates high-frequency information of the image and is high in variance. The signal variance decreases with inversion time, and the predicted noise’s variance increases with inversion time. The effective DDIM Inversion(Song, Meng, and Ermon 2020) iteration is:
$\mathbf{x}_{t+1}=$ $\underbrace{\left(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}}_{\text{signal / mean},\,\mathbf{z}^{\text{inv}}_{\mu,t+1}}$ (2)
$\hskip 40.00006pt+\underbrace{\sqrt{1-\alpha_{t+1}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)}_{\text{noise / variance},\,\mathbf{z}^{\text{inv}}_{\sigma,t+1}},$
where $\mathbf{x}_{t+1}$ is the noisy latent at timestep $t+1$. $\mathbf{x}_{t}$ is the latent at timestep $t$, which is less noisy than $\mathbf{x}_{t+1}$. $\alpha_{t}$ and $\alpha_{t+1}$ are parameters controlling the noise schedule at timesteps $t$ and $t+1$, derived from the noise variance schedule. $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ is the predicted noise at timestep $t$, estimated by the diffusion U-Net during the reverse process.
Rather than inverting all the way to $t=T$, where the latent is similar to white noise and the reverse trajectory becomes unstable, we stop at an intermediate timestep $t^{*}<T$. At $t^{*}$, the DDIM latent still preserves enough low‑frequency structure to support direct view translation via our TUNet.
The signal/mean in Equation (2) is the coarse-grained image representation on which we perform the view translation. In addition, the noise/variance in Equation (2) encodes image-specific features that are recovered during the denoising process (Staniszewski, Kuciński, and Deja 2025). For the task of novel-view synthesis, this noise/variance can be used to induce high-frequency details into the view-transformed latent, which can then be fed to DDIM sampling. Based on the aforementioned discussion, we formalize two things for the task of novel view synthesis:
  • Spectral bias of the diffusion model can be exploited to perform the view transformation in the low-frequency space with our translation network, TUnet.
  • To compensate for high frequency details, we utilize the noise/variance term of DDIM inversion in Equation (2) to formulate a fusion strategy.
DDIM-inverted Latents
Let $\mathbf{z}^{\text{inv}}_{t}$ be the DDIM-inverted latent. If we use this latent at $t=T$, we may see that the DDIM sampled image deviates from the input image, especially when we do it in fewer DDIM steps (Bao et al. 2025; Zeng, Suganuma, and Okatani 2025; Feng et al. 2024). Thus, we fix $t$ = 600 and get our DDIM inverted noisy initial latent in 30 DDIM steps. Further, the signal/mean term $\mathbf{z}^{\text{inv}}_{\mu,t+1}$ of Equation (2) is the diffusion network’s estimate of the clean latent which we obtain at $t$ by denoising $\mathbf{z}^{\text{inv}}_{t}$ according to the diffusion score model $\boldsymbol{\epsilon}_{\theta}$. This signal/mean term is what we feed into TUNet for view translation. We visualise the reconstructed image corresponding to signal/mean term in Figure 3. We observe that the reconstructed image primarily comprises of the low‑frequency components of the input image. Therefore, in order to impose high-frequency information, we make use of the noise/variance term $\mathbf{z}^{\text{inv}}_{\sigma,t+1}$ from Equation (2) that re‑injects the predicted noise at the level $t$. We utilize a pretrained latent diffusion model (LDM) (Rombach et al. 2022) as our generative prior, which we use to compute the DDIM-inverted latents. We first describe the TUNet model, followed by the Fusion Strategy. As we fix timestep $t$ at 600, we drop the subscript $t$ while representing signal/mean and noise/variance terms: $\mathbf{z}^{\text{inv}}_{\mu,t+1}$ and $\mathbf{z}^{\text{inv}}_{\sigma,t+1}$ in Equation (2) from now on.
Figure 3: Left: Mean of the DDIM inverted latent at tt=600. Latent is decoded using VAE for visualization. Right: Original $\times512$ image.
TUNET Architecture
TUNet is a U-Net  (Ronneberger, Fischer, and Brox 2015) inspired encoder-decoder architecture designed to predict the DDIM-inverted latent’s mean representation of the target view. TUNet introduces cross attention between an input or reference view and a target view, enabling effective feature transfer between viewpoints. The architecture is conditioned on both camera parameters and class embeddings at multiple stages to preserve geometric consistency and semantic integrity.
Input and Conditioning
The input image $\mathbf{x_{\text{ref}}}$ is initially mapped to the latent space using a VAE encoder, yielding $\mathbf{z_{\text{ref}}}$. We then perform DDIM inversion on this latent space representation $\mathbf{z_{\text{ref}}}$ to obtain the mean term $\mathbf{z^{\text{inv}}_{\text{ref},\mu}}$, which acts as input to TUNet. The following information is used as a condition to TUNet at various stages:
  • Camera Embedding $\mathbf{C}$ = $\mathbf{(K,R,t)}$: A vectorized form of camera intrinsics $\mathbf{K}$ and extrinsics $\mathbf{(R,t)}$ is passed through a learnable linear layer to produce an embedding vector $\mathbf{e}_{C}\in\mathbb{R}^{d_{C}}$.
  • Class Embedding: A learnable class embedding corresponding to the scene category, mapped to $\mathbf{e}_{c}\in\mathbb{R}^{d_{c}}$.
These embeddings are concatenated with the time embedding $\boldsymbol{\gamma}(t)$ $\in$ $\mathbb{{R}}^{d_{t}}$, and the combined vector $\bigl{[}\boldsymbol{\gamma}(t)\oplus\mathbf{e}_{C}\oplus\mathbf{e}_{c}\bigr{]}$} $\in$ $\mathbb{R}^{d_{t}+d_{C}+d_{c}}$ is passed through a learnable linear projection to align it with the time embedding space. The resulting projected vector is broadcast spatially and added to the feature maps $\mathbf{f}$ at each downsampling, mid, and upsampling block:
$\mathbf{f}^{\prime}=\mathbf{f}+{\text{Proj}}_{{\text{combined}}}{\bigl{[}\boldsymbol{\gamma}(t)\oplus\mathbf{e}_{C}\oplus\mathbf{e}_{c}\bigr{]}},$
where $\text{Proj}_{\text{combined}}$} is a learned linear layer mapping the concatenated embedding to $\mathbb{R}^{d_{t}}$. This enables joint conditioning on time, camera viewpoint, and scene class.
Encoder (Down Blocks)
The encoder comprises a series of residual downsampling blocks that reduce spatial resolution while expanding the depth of the feature. Each block is conditioned on the camera and the class embeddings of the input or reference view ($\mathbf{C_{\text{ref}},c_{\text{ref}}}$). These embeddings are added after concatenation and projection:
$\mathbf{f}^{(i)}=\text{Down}_{i}\bigl{(}\mathbf{f}^{(i-1)}+\text{Proj}_{\mathrm{combined}}\bigl{[}\boldsymbol{\gamma}(t)\oplus\mathbf{e}_{\mathbf{C}_{\mathrm{ref}}}\oplus\mathbf{e}_{\mathbf{c}_{\mathrm{ref}}}\bigr{]}\bigr{)},$
where $i$ denotes the depth of the block in TUNet.
Bottleneck and Decoder (Mid + Up Blocks)
The bottleneck block is conditioned on both the input or reference and target view camera embeddings ($\mathbf{C_{\text{ref}}}$, $\mathbf{C_{\text{tar}}}$) along with the class embeddings, allowing the model to capture viewpoint transitions at the latent level. The upsampling stages are conditioned only on the target view’s camera and class embeddings ($\mathbf{C_{\text{tar}}}$, $\mathbf{c_{\text{tar}}}$), guiding the representation toward the desired target view:
$\mathbf{f}^{\mathrm{mid}}=\text{Mid}\bigl{(}\mathbf{f}^{\mathrm{enc}}+\text{Proj}_{\mathrm{combined}}\bigl{[}\boldsymbol{\gamma}(t)\oplus\mathbf{e}_{{C}_{\text{ref}}}\oplus\mathbf{e}_{{C}_{\text{tar}}}\oplus\mathbf{e}_{{c}_{\text{tar}}}\bigr{]}\bigr{)},$
$\mathbf{f}^{(i)}=\text{Up}_{i}\bigl{(}\mathbf{f}^{(i-1)}+\text{Proj}_{\mathrm{combined}}\bigl{[}\boldsymbol{\gamma}(t)\oplus\mathbf{e}_{{C}_{\mathrm{tar}}}\oplus\mathbf{e}_{{c}_{\mathrm{tar}}}\bigr{]}\bigr{)}.$
Cross-Attention Module
A cross-attention mechanism is integrated in the mid and up blocks, enabling information flow from the reference to the target view using ray information and latent feature alignment. Let $\mathbf{r_{\text{ref}}}$ denote the ray embeddings of the reference view and $\mathbf{r_{\text{tar}}}$ denote the ray embeddings of the target view. We use standard ray parameterization as in NeRF (Mildenhall et al. 2021) to compute ray origins and directions for camera pose encoding to get $\mathbf{r_{\text{ref}}}$ and $\mathbf{r_{\text{tar}}}$. Let $\mathbf{z^{\text{inv}}_{\text{ref},\mu}}$ be the DDIM-inverted latent mean of the reference image, and $\mathbf{f_{\text{tar}}}$ be the intermediate target feature maps at the cross-attention block.
The attention mechanism uses the formulation:
$\mathbf{Q}=\mathbf{W}_{Q}[\mathbf{r}_{\text{tar}}\mathbin{\|}\mathbf{f}_{\text{tar}}],\ \mathbf{K}=\mathbf{W}_{K}[\mathbf{r}_{\text{ref}}\mathbin{\|}\mathbf{z}^{\text{inv}}_{\text{ref},\mu}],\ \mathbf{V}=\mathbf{W}_{V}\mathbf{z}^{\text{inv}}_{\text{ref},\mu}$
${{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})}={softmax\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)\mathbf{V}}.$ (3)
The output of attention is then added back to the target features:
$\mathbf{f}^{\prime}_{\mathbf{\text{tar}}}\;=\;\mathbf{f}_{\mathbf{\text{tar}}}\;+\;{Attn}\bigl{(}\mathbf{Q},\,\mathbf{K},\,\mathbf{V}\bigr{)}.$
The output of TUNet is a latent $\tilde{\mathbf{z}}^{\text{inv}}_{\text{tar},\mu}$ representing the synthesized view’s DDIM inverted mean term corresponding to the target camera. Using $\tilde{\mathbf{z}}^{\text{inv}}_{\text{tar},\mu}$, we next explain the fusion strategy.
Fusion Strategy
To synthesize semantically rich target view latents from the predicted DDIM-inverted mean latent $\tilde{\mathbf{z}}^{\text{inv}}_{\text{tar},\mu}$, we introduce two fusion strategies that combine this mean latent with a noise component derived from the input view latent. These strategies re‑inject the learned noise variance, that is, the high‑frequency details, into the coarse latent. We utilize the fact that the noise/variance term Equation (2) in the DDIM-inverted latent of the input view contains scene-level attributes and characteristics (Staniszewski, Kuciński, and Deja 2025), which can be used to synthesize the scene from a novel view when fused with TUNet’s prediction.
Strategy A: Variance Fusion via $\sigma$-Component
In this strategy, we explicitly extract the variance (or noise) component from the DDIM-inverted latent of the input view, denoted as $\mathbf{z}_{\text{ref},\sigma}^{\text{inv}}$. We perform DDIM inversion on $\mathbf{z}_{\text{ref}}$, and extract the equivalent noise/variance term $\mathbf{z}_{\text{ref},\sigma}^{\text{inv}}$ from Equation (2). The final latent is computed as:
$\mathbf{z}_{\text{noisy}}=\tilde{\mathbf{z}}_{\text{tar},\mu}^{\text{inv}}+\mathbf{z}_{\text{ref},\sigma}^{\text{inv}}.$ (4)
The fused latent $\mathbf{z}_{\text{noisy}}$ is then passed into the Stable Diffusion U-Net to compute the noise prediction:
$\boldsymbol{\epsilon}_{\theta}=\text{U-Net}(\mathbf{z}_{\text{noisy}},t).$ (5)
The initial latent for DDIM sampling is obtained as:
$\tilde{\mathbf{z}}_{\text{tar}}^{\text{inv}}=\tilde{\mathbf{z}}_{\text{tar},\mu}^{\text{inv}}+\sqrt{1+\alpha_{t+1}}\boldsymbol{\epsilon}_{\theta}.$ (6)
Strategy B: Direct Noise Addition from Reference Inversion
Here, we directly use the noise component from the full DDIM-inverted latent of the input view $\mathbf{z}_{\text{ref}}^{\text{inv}}$, rather than extracting its variance separately. The initial latent $\tilde{\mathbf{z}}_{\text{tar}}^{\text{inv}}$ for DDIM sampling is computed as :
$\tilde{\mathbf{z}}_{\text{tar}}^{\text{inv}}=\tilde{\mathbf{z}}_{\text{tar},\mu}^{\text{inv}}+\sqrt{1+\alpha_{t+1}}\mathbf{z}_{\text{ref}}^{\text{inv}}.$ (7)
We generate samples using $\tilde{\mathbf{z}}_{\text{tar}}^{\text{inv}}$ from both Equation (6) and Equation (7), and pick the best generated result where we pick the result which has lower LPIPS score.
Figure 4: Qualitative results with our 167-class trained model.
Training Objective
Our training objective is to align the DDIM-inverted latent mean of the prediction and ground-truth. We achieve this by minimizing the Mean Squared Error (MSE) loss between the predicted target latent mean $\tilde{\mathbf{z}}^{\text{inv}}_{\text{tar},\mu}$ and the ground-truth DDIM-inverted latent mean of the target view $\mathbf{z}_{\text{tar},\mu}^{\text{inv}}$:
$\mathcal{L}_{\text{MSE}}=\left\|\tilde{\mathbf{z}}^{\text{inv}}_{\text{tar},\mu}-\mathbf{z}_{\text{tar},\mu}^{\text{inv}}\right\|_{2}^{2}.$ (8)
Experiments
Dataset: MvImgNet (Yu et al. 2023b) consists of 6.5 million frames of real-world scenes across 238 categories. We use two subsets of MVImgNet (i) three scene categories: sofas, chairs, and tables. A 90-5-5 (Anciukevičius et al. 2024) split is used for training, validation, and testing, respectively, determined by lexicographic ordering of the scene identifiers. (ii) We use 8.5 lakh frames across 167 classes and for each class, we keep 1 scene out of 99 in the test set to evaluate our results and compare with other methods.
Pre-Processing: We resize the shorter dimension of the images to be 512 and resize the other dimension to maintain the aspect ratio and then take centre crop of $512\times 512$. These $512\times 512$ RGB images are subsequently passed through VAE encoder and the DDIM inversion pipeline to get the inverted latents $\mathbf{z^{\text{inv}}}$ and extract their mean and variance components. We perform DDIM inversion from $t$ = 0 till $t$ = 600 in 30 steps. The data in the inverted latent space $\mathbf{z^{\text{inv}}}$ is of dimension $4\times 64\times 64$.
Implementation Details : TUnet has approximately 148M parameters. The dimensions of both class and camera embeddings are 64, and the cross-attention dimension is 768 with an attention head dimension of 64. We use the latent diffusion backbone (Rombach et al. 2022). For training, we randomly pair frames 1-10 of each scene with frames 15-25. We effectively use 20 frames per scene for training. We adopt the same frame pairing strategy for our evaluation. We train two models on our subset (i) 3 classes and (ii) 167 classes. Our 167 class model is trained for 450 epochs on a single 40GB A400 for 17 GPU days with a batch size of 32 and a learning rate of 1e-5, and we decay the learning rate using a cycle scheduler. During inference, we generate final results with 30 DDIM sampling steps with the initial latent being Equation (6) or Equation (7), which represents the noisy latent at $t$ = 600.
We compare our 3-class model with GIBR (Anciukevičius et al. 2024) and the 167-class model with NViST (Jang and Agapito 2024). For GIBR, we have the same train/test split and directly report the results from their paper. For NViST, we use their pre-trained model to test exact input/target frame pairs. All of the testing frames are from unseen scenes within the classes used in training. To compare with GIBR, we resize our $512\times 512$ results to $256\times 256$. For direct comparison with NViST, we resize our results to $90\times 90$. We report LPIPS and FID scores with the resized results for comparisons. We report PSNR and SSIM with the original $512\times 512$ resolution results.
Quantitative Comparison
3-class model: Comparison with GIBR on 3 classes at a resolution of $256\times 256$ is shown in Table 1. Input and target pairs from 168 unseen scenes are used for testing. We outperform GIBR in terms of LPIPS. GIBR trains the entire diffusion process in the RGB space and also uses multiple views while training and volume rendering to generate the final image. Thus, GIBR does better in terms of PSNR and SSIM. However, training diffusion in pixel space is very expensive. On the other hand, we only train our TUNet with 148M parameters using latents.
Method LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
GIBR $0.510$ 17.61 0.554
Ours 0.490 $15.71$ $0.523$
Table 1: Comparison for 3 classes - chairs, sofa, tables. Resolution is $256\times 256$. (Note: Exact setting of GIBR is not reproducible as the code is not available.)
167-class model: Comparison with NViST at a resolution of $90\times 90$ is shown in Table 2. Input and target pairs from 360 unseen scenes are used for testing. Here, we see that our method performs better in terms of LPIPS, PSNR, SSIM, and FID.
Method LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$
NViST 0.448{0.448} 14.31{14.31} $0.566$ $91.63$
Ours 0.409 16.16 0.578 65.50
Table 2: Comparison for 167 classes. Resolution is $90\times 90$
We show the synthesized results in Figure 4. In the case of the kettle, we can see that the unobserved region is synthesized with high fidelity. Similarly, in the case of a bowl (third row, middle column), the shadow is faithfully synthesized.
Qualitative Comparison
We compare our results with NViST (Jang and Agapito 2024) in Figure 5. It is evident that our method synthesizes the target with high fidelity and is able to generate results for near as well as far target views, where NViST fails.
Figure 5: We resize our results to $\times90$ to show comparison with NViST on unseen test scenes from 5 classes.
Unseen classes: We show qualitative results on 6 unseen classes in Figure 6. We cover outdoor and indoor scenes, as well as include large and small object classes in the test set. Even on unseen classes, we are able to generate high-resolution reconstruction for a diverse set of scenes.
Figure 6: Results on 6 unseen classes from MVImgNet
Out of domain data: To evaluate zero‑shot generalization beyond MVImgNet, we assembled an out‑of‑domain test set by downloading freely‑licensed photographs from Unsplash (Unsplash Contributors 2025) featuring natural scenes. We show the results and compare with Zero123++ (Shi et al. 2023) in Figure 7. While both methods successfully generate plausible novel viewpoints, our approach produces more faithful surface textures and preserves natural scene characteristics.
Figure 7: Out-of-domain images
Ablation
The ablation results are presented in Table 3. We evaluate the following settings. In the first setting, Concat, we concatenate class and camera embeddings with the input, but do not inject them into every ResNet block. Here, we see that there is a significant performance drop. Second, w/o cross-attn, where we remove all cross-attention layers. The results degrade for all three metrics.
Setting LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
Concat (class, cam. w inp.) 0.508 15.38 0.516
w/o cross-attention 0.506 15.41 0.515
Full model 0.492 15.71 0.523
Table 3: Ablation study using 3-class.
Conclusion
In this work, we propose a novel method using TUNet and a fusion strategy to synthesize high-quality novel views. Our method synthesizes the novel views using single input image and camera parameters. Compared to prior works, which train a heavy diffusion model, our method trains a lightweight translation network to obtain view translation in latent space. To enrich the predicted latent with high frequency scene details, we propose a novel fusion strategy. Our experiments reveal strong performance under various settings.
Supplementary Material
Novel View Synthesis using DDIM Inversion
Supplementary Materials
Novel View Synthesis and Editing
As we work with DDIM inverted noise, we have the flexibility to extend our pipeline to editing via textual prompts during the final sampling. For our novel view reconstruction, we use BLIP (Li et al. 2022) generated captions for the input image $\mathbf{x}_{ref}$. We edit these prompts to add desired scenarios in our reconstructed novel view. Figure 8 shows results of our editing mechanism.
Figure 8: We use textual prompts to edit the novel view reconstructions as desired. In the first row, we place a bag on the bench. In the second row, a dog is introduced in the scene. In the third row, a pair of shoes are placed by the bench.
Synthesis of Multiple Views from Single Image
In Figure 9, we generate multiple frames using a single input image. We query the same input image with multiple target camera parameters and reconstruct the novel views with respect to different target views. Even for long-range viewpoints, the proposed method achieves good synthesis results.
Figure 9: Generating multiple frames with single input image.
Qualitative Ablation Results
In Figure 10, we show qualitative results with our ablation setting. We can observe that in the ‘w/o Cross Attention’ setting, the view transformation geometry suffers. In the ‘Concat’ setting, the object loses fine-grained details while synthesizing the novel view. Full Model shows the best performance.
Figure 10: Ablation
Additional Results
In Table 4, we report the results on the original synthesized resolution of $512\times 512$.
Method LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$
Ours $0.556$ $16.162$ $0.578$ $69.604$
Table 4: Our results at original resolution of $512\times 512$
Figure 11: Qualitative results with our 3-class trained model on MVImgNet test set.
Figure 12: Additional Qualitative results with our 167-class trained model on MVImgNet test set.
We show additional results using our 3-class and 167-class models in Figures 11 and 12, respectively.

### v2

@@
 To synthesize the novel view, we use the fused latent as the initial condition for DDIM sampling, leveraging the generative prior of the pretrained diffusion model.
-Extensive experiments on MVImgNet demonstrate that our method outperforms existing methods.
-Code is available.
+Extensive experiments on MVImgNet and RealEstate10K demonstrate that our method outperforms existing methods.
+The code is available at https://github.com/Visual-Conception-Group/ddim_nvs .
 (a) MvImgNet test set
@@
 Figure 1: (a) High-resolution ($512\times 512$) novel-view synthesis on the MvImgNet test set from a single input image and camera parameters, (b) Zero-shot synthesis on out-of-domain images downloaded from Unsplash.
-Introduction
+1 Introduction
 Novel view synthesis is a fundamental task in computer vision and graphics.
-Remarkable works such as NeRFs (Mildenhall et al.
-2021) and 3DGS (Kerbl et al.
-2023) are extensively used in 3d scene understanding.
+Remarkable works such as NeRFs [30] and 3DGS [23] are extensively used in 3d scene understanding.
 Several works improve upon these foundational works.
 However, their dependence on scene-level optimization and the need for a dense set of views limit usability.
-Diffusion models (Rombach et al.
-2022; Podell et al.
-2023), have gained significant traction for the task of novel view synthesis (Elata et al.
-2025; Tang et al.
-2025).
-A classical approach is to fine-tune these models on 3D datasets along with a module that encodes the 3D geometry into the architecture (Tang et al.
-2023; Liu et al.
-2024; Long et al.
-2024; Huang et al.
-2024; Bourigault and Bourigault 2024; Gao et al.
-2024).
+Diffusion models [35, 33], have gained significant traction for the task of novel view synthesis [10, 44].
+A classical approach is to fine-tune these models on 3D datasets along with a module that encodes the 3D geometry into the architecture [43, 27, 29, 21, 3, 14].
 However, the generated outputs lack consistency in multiview reconstruction as the generation is not entirely controllable and results in images of inadequate quality, and often creates blurry results for long-range viewpoint reconstruction.
-DDIM (Song, Meng, and Ermon 2020) proposed a deterministic inversion “DDIM Inversion”, which sequentially adds noise to an image to obtain a noisy latent.
+DDIM [40] proposed a deterministic inversion “DDIM Inversion”, which sequentially adds noise to an image to obtain a noisy latent.
 The noisy latent can be retraced to the original image using DDIM sampling.
 This latent encapsulates the signal and the noise that contribute to the mean and variance, which changes the distribution of the noise latent at each inversion time.
-Previous works such as (Garibi et al.
-2024; Mokady et al.
-2023) try to optimize or configure this noise representation to better align with the given task.
-(Staniszewski, Kuciński, and Deja 2025) study inversion noise in detail and claim that the DDIM inversion latent space is less manipulative, which makes direct interpolation with this noise latent difficult for tasks such as novel view synthesis and editing.
+Previous works such as [15, 31] try to optimize or configure this noise representation to better align with the given task.
+[41] study inversion noise in detail and claim that the DDIM inversion latent space is less manipulative, which makes direct interpolation with this noise latent difficult for tasks such as novel view synthesis and editing.
 This paper proposes a method to generate a novel view from a given input image and camera parameters.
@@
 This mapping only approximates a coarse-grained version of the target view.
-This is due to the fact that diffusion models exhibit spectral bias and favor low-frequency components (Choi et al.
-2022).
+This is due to the fact that diffusion models exhibit spectral bias and favor low-frequency components [7].
 In order to induce the high frequency components, we introduce a novel noisy latent fusion strategy.
@@
   • In our experiments, we show that the method achieves superior results in terms of LPIPS, PSNR, SSIM, and FID.
-Related Work
-Neural Radiance Field: Neural field approaches, such as Neural Radiance Fields (NeRF) (Mildenhall et al.
-2021), use learnable functions to map 3D spatial coordinates and viewing directions to volumetric density and color.
+2 Related Work
+Neural Radiance Field: Neural field approaches, such as Neural Radiance Fields (NeRF) [30], use learnable functions to map 3D spatial coordinates and viewing directions to volumetric density and color.
 These models synthesize novel views by performing volumetric rendering via ray marching through the learned scene representation.
 NeRF has demonstrated that high-quality novel views can be rendered when trained on a dense set of input views.
-While recent extensions such as PixelNeRF (Yu et al.
-2021), IBRNet (Wang et al.
-2021), MultiDiff (Müller et al.
-2024), and others (Henzler et al.
-2021; Liu et al.
-2022; Wu et al.
-2023) aim to perform view synthesis from fewer input views, they often suffer in regions with missing or occluded content.
+While recent extensions such as PixelNeRF [59], IBRNet [49], MultiDiff [müller2024multidiffconsistentnovelview], and others [18, 28, 54] aim to perform view synthesis from fewer input views, they often suffer in regions with missing or occluded content.
 Because these models make deterministic predictions without explicit uncertainty modeling, the generated output tends to average over ambiguities, leading to blurry and less plausible reconstruction in unobserved regions.
-Gaussian Splatting: 3D Gaussian Splatting (3DGS) (Kerbl et al.
-2023; Peng et al.
-2024; Zheng et al.
-2025) represent scenes using a set of anisotropic 3D Gaussians.
-Gaussian Splatting methods are deterministic and depend heavily on accurate multi-view geometry or densely sampled camera poses (Li et al.
-2024).
+Gaussian Splatting: 3D Gaussian Splatting (3DGS) [23, 32, 65] represent scenes using a set of anisotropic 3D Gaussians.
+Gaussian Splatting methods are deterministic and depend heavily on accurate multi-view geometry or densely sampled camera poses [26].
 When applied in sparse-view or single-view settings, they often fail to generate plausible content in unseen regions because they lack generative priors.
@@
 This setting spans both short and long-range viewpoint changes.
-Under such conditions, methods such as NeRF (Mildenhall et al.
-2021) and 3DGS (Kerbl et al.
-2023) struggle to extrapolate effectively from a single image, even when augmented with generative guidance as in (Szymanowicz et al.
-2024; Tewari et al.
-2023).
-Transformers: NViST (Jang and Agapito 2024) adopts a transformer-based encoder-decoder architecture (Vaswani et al.
-2017; Dosovitskiy et al.
-2021) to predict a radiance field from a single image, enabling novel view synthesis via NeRF-style volumetric rendering.
+Under such conditions, methods such as NeRF [30] and 3DGS [23] struggle to extrapolate effectively from a single image, even when augmented with generative guidance as in [42, 45].
+Transformers: GeoGPT [36] was one of the early works to perform view synthesis using transformers.
+NViST [22] adopts a transformer-based encoder-decoder architecture [48, 9] to predict a radiance field from a single image, enabling novel view synthesis via NeRF-style volumetric rendering.
 However, NViST suffers from loss of fine details due to aggressive downsampling (by a factor of 12), and it struggles to synthesize long-range viewpoints (when the target frame is more than 15 frames away from the input frame in a 30‑frame sweep).
@@
 In the following, we identify them as the ones which finetune pre-trained diffusion models, or train diffusion models from scratch.
-Pretrained Diffusion Models: MVDiffusion (Tang et al.
-2023), Zero123++ (Shi et al.
-2023), SyncDreamer (Liu et al.
-2024), Wonder3D (Long et al.
-2024), EpiDiff (Huang et al.
-2024), BoostDream (Yu et al.
-2024), MVDiff (Bourigault and Bourigault 2024), CAT3D (Gao et al.
-2024), Magic-Boost (Yang et al.
-2024), Cycle3D (Tang et al.
-2025), use a pretrained or finetuned diffusion model.
+Pretrained Diffusion Models: MVDiffusion [43], Zero123++ [39], SyncDreamer [27], Wonder3D [29], EpiDiff [21], BoostDream [62], MVDiff [3], CAT3D [14], Magic-Boost [57], Cycle3D [44], GenWarp [38], use a pretrained or finetuned diffusion model.
 MVDiffusion modifies the Stable Diffusion architecture by introducing a cross-branch attention mechanism, known as correspondence-aware attention (CAA), to model inter-view dependencies.
 SyncDreamer constructs a view frustum feature volume from all the target noisy views and injects these into the pretrained denoising Unet using depth-wise attention layers.
-EpiDiff (Huang et al.
-2024) integrates an attention module guided by epipolar constraints into the intermediate and decoding stages of the U-Net, enabling the model to capture generalized epipolar geometry across views.
+EpiDiff [21] integrates an attention module guided by epipolar constraints into the intermediate and decoding stages of the U-Net, enabling the model to capture generalized epipolar geometry across views.
+GenWarp [38] introduces a warp and inpaint technique.
 Most existing approaches either fine-tune the diffusion model or inject spatial features corresponding to the target view into the base model’s denoising U-Net.
@@
 Instead, we provide external conditioning input to TUNet to obtain the latent corresponding to the target view.
-The latent text space of diffusion models has also been explored for view synthesis Burgess, Wang, and Yeung-Levy (2024).
-However, it is typically limited to small-scale datasets and is not directly applicable to scene-level novel view synthesis tasks.
 Figure 2: Overview: Given a single reference image $\mathbf{x_{\text{ref}}}$, we first apply DDIM inversion up to $t=600$ to obtain the mean latent $\mathbf{z}_{\text{ref},\mu}^{\text{inv}}$.
@@
 Training Diffusion Model from Scratch: Several recent works train diffusion models from scratch for novel view synthesis, including Tseng et al.
-(2023), Photometric-NVS (Yu et al.
-2023a), DiffDreamer (Cai et al.
-2023), GIBR (Anciukevičius et al.
-2024), and (Henderson et al.
-2024).
-Photometric-NVS (Yu et al.
-2023a) introduces a two-stream latent diffusion architecture that independently processes the source and noisy target views, while exchanging information via pose-conditioned cross-attention mechanisms.
-GIBR (Anciukevičius et al.
-2024) models 3D scenes using IB-planes and trains the diffusion process directly in pixel space, enabling learning of a joint distribution over multi-view observations and camera poses.
+[46], Photometric-NVS [60], DiffDreamer [5], GIBR [1], and [17].
+Photometric-NVS [60] introduces a two-stream latent diffusion architecture that independently processes the source and noisy target views, while exchanging information via pose-conditioned cross-attention mechanisms.
+GIBR [1] models 3D scenes using IB-planes and trains the diffusion process directly in pixel space, enabling learning of a joint distribution over multi-view observations and camera poses.
 Training entire diffusion models end-to-end is computationally expensive and requires large-scale datasets to achieve high-resolution and photorealistic reconstruction.
@@
 The final novel view is synthesized using a pretrained diffusion pipeline, which decodes the predicted latent.
-Method
+3 Method
 Given a single reference image and camera parameters of the target viewpoint, our work addresses the task of novel view synthesis.
@@
 Our method is illustrated in Figure 2.
-Spectral Behavior of Diffusion
-In (Kingma and Gao 2023; Falck et al. 2025), authors study the spectral behavior of diffusion. The forward diffusion (Ho, Jain, and Abbeel 2020) process is given by:
+3.1 Spectral Behavior of Diffusion
+In [24, 11], authors study the spectral behavior of diffusion. The forward diffusion [19] process is given by:
 $\mathbf{x}_{t}=\underbrace{\sqrt{\overline{\alpha}_{t}}\,\mathbf{x}_{0}}_{\text{signal}}+\underbrace{\sqrt{1-\overline{\alpha}_{t}}\,\boldsymbol{\epsilon}}_{\text{noise}},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
 where $\mathbf{x}_{0}$ is the clean latent, $\mathbf{x}_{t}$ is the latent corresponding to the timestep $t$ and $\bar{\alpha}_{t}$ is the scaling factor.
-High-frequency components, representing fine details, are degraded more rapidly and prior to low-frequency components during the forward diffusion process (Kingma and Gao 2023; Falck et al.
-2025).
+High-frequency components, representing fine details, are degraded more rapidly and prior to low-frequency components during the forward diffusion process [24, 11].
 This property is also consistent in the reverse process of diffusion.
 As shown in Choi et al.
-(2022), diffusion models inherently favor lower frequencies, which implies that more emphasis must be placed on modelling high-frequency details.
-In addition, the noise component is often observed to deviate from a standard multivariate Gaussian distribution (Staniszewski, Kuciński, and Deja 2025).
+[7], diffusion models inherently favor lower frequencies, which implies that more emphasis must be placed on modelling high-frequency details.
+In addition, the noise component is often observed to deviate from a standard multivariate Gaussian distribution [41].
 At later iterations of inversion, the noise encapsulates high-frequency information of the image and is high in variance.
 The signal variance decreases with inversion time, and the predicted noise’s variance increases with inversion time.
-The effective DDIM Inversion(Song, Meng, and Ermon 2020) iteration is:
-$\mathbf{x}_{t+1}=$ $\underbrace{\left(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}}_{\text{signal / mean},\,\mathbf{z}^{\text{inv}}_{\mu,t+1}}$ (2)
-$\hskip 40.00006pt+\underbrace{\sqrt{1-\alpha_{t+1}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)}_{\text{noise / variance},\,\mathbf{z}^{\text{inv}}_{\sigma,t+1}},$
+The effective DDIM Inversion[40] iteration is:
+$\mathbf{x}_{t+1}=$ $\underbrace{\left(\mathbf{x}_{t}-\sqrt{1-\overline{\alpha}_{t}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)\sqrt{\frac{\overline{\alpha}_{t+1}}{\overline{\alpha}_{t}}}}_{\text{signal / mean},\,\mathbf{z}^{\text{inv}}_{\mu,t+1}}$ (2)
+$\hskip 40.00006pt+\underbrace{\sqrt{1-\overline{\alpha}_{t+1}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)}_{\text{noise / variance},\,\mathbf{z}^{\text{inv}}_{\sigma,t+1}},$
 where $\mathbf{x}_{t+1}$ is the noisy latent at timestep $t+1$.
-$\mathbf{x}_{t}$ is the latent at timestep $t$, which is less noisy than $\mathbf{x}_{t+1}$.
-$\alpha_{t}$ and $\alpha_{t+1}$ are parameters controlling the noise schedule at timesteps $t$ and $t+1$, derived from the noise variance schedule.
 $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ is the predicted noise at timestep $t$, estimated by the diffusion U-Net during the reverse process.
@@
 The signal/mean in Equation (2) is the coarse-grained image representation on which we perform the view translation.
-In addition, the noise/variance in Equation (2) encodes image-specific features that are recovered during the denoising process (Staniszewski, Kuciński, and Deja 2025).
+In addition, the noise/variance in Equation (2) encodes image-specific features that are recovered during the denoising process [41].
 For the task of novel-view synthesis, this noise/variance can be used to induce high-frequency details into the view-transformed latent, which can then be fed to DDIM sampling.
@@
   • To compensate for high frequency details, we utilize the noise/variance term of DDIM inversion in Equation (2) to formulate a fusion strategy.
-DDIM-inverted Latents
+3.2 DDIM-inverted Latents
 Let $\mathbf{z}^{\text{inv}}_{t}$ be the DDIM-inverted latent.
-If we use this latent at $t=T$, we may see that the DDIM sampled image deviates from the input image, especially when we do it in fewer DDIM steps (Bao et al.
-2025; Zeng, Suganuma, and Okatani 2025; Feng et al.
-2024).
+If we use this latent at $t=T$, we may see that the DDIM sampled image deviates from the input image, especially when we do it in fewer DDIM steps [2, 63, 12].
 Thus, we fix $t$ = 600 and get our DDIM inverted noisy initial latent in 30 DDIM steps.
@@
 Therefore, in order to impose high-frequency information, we make use of the noise/variance term $\mathbf{z}^{\text{inv}}_{\sigma,t+1}$ from Equation (2) that re‑injects the predicted noise at the level $t$.
-We utilize a pretrained latent diffusion model (LDM) (Rombach et al.
-2022) as our generative prior, which we use to compute the DDIM-inverted latents.
+We utilize a pretrained latent diffusion model (LDM) [35] as our generative prior, which we use to compute the DDIM-inverted latents.
 We first describe the TUNet model, followed by the Fusion Strategy.
 As we fix timestep $t$ at 600, we drop the subscript $t$ while representing signal/mean and noise/variance terms: $\mathbf{z}^{\text{inv}}_{\mu,t+1}$ and $\mathbf{z}^{\text{inv}}_{\sigma,t+1}$ in Equation (2) from now on.
-Figure 3: Left: Mean of the DDIM inverted latent at tt=600. Latent is decoded using VAE for visualization. Right: Original $\times512$ image.
-TUNET Architecture
-TUNet is a U-Net  (Ronneberger, Fischer, and Brox 2015) inspired encoder-decoder architecture designed to predict the DDIM-inverted latent’s mean representation of the target view.
+Figure 3: Mean of the DDIM inverted latent at $t=400,600,800$, respectively.
+Latent is decoded using VAE for visualization.
+Original $\times512$ image.
+At $t=400$, the mean reflects dominant low frequencies which precludes generation of diverse images.
+At $t=800$, the low frequency component is extremely weak.
+$t=600$ provides a weak yet effective signal for translation.
+3.3 TUNET Architecture
+TUNet is a U-Net  [37] inspired encoder-decoder architecture designed to predict the DDIM-inverted latent’s mean representation of the target view.
 TUNet introduces cross attention between an input or reference view and a target view, enabling effective feature transfer between viewpoints.
 The architecture is conditioned on both camera parameters and class embeddings at multiple stages to preserve geometric consistency and semantic integrity.
-Input and Conditioning
+3.3.1 Input and Conditioning
 The input image $\mathbf{x_{\text{ref}}}$ is initially mapped to the latent space using a VAE encoder, yielding $\mathbf{z_{\text{ref}}}$.
@@
 where $\text{Proj}_{\text{combined}}$} is a learned linear layer mapping the concatenated embedding to $\mathbb{R}^{d_{t}}$. This enables joint conditioning on time, camera viewpoint, and scene class.
-Encoder (Down Blocks)
+3.3.2 Encoder (Down Blocks)
 The encoder comprises a series of residual downsampling blocks that reduce spatial resolution while expanding the depth of the feature.
@@
 where $i$ denotes the depth of the block in TUNet.
-Bottleneck and Decoder (Mid + Up Blocks)
+3.3.3 Bottleneck and Decoder (Mid + Up Blocks)
 The bottleneck block is conditioned on both the input or reference and target view camera embeddings ($\mathbf{C_{\text{ref}}}$, $\mathbf{C_{\text{tar}}}$) along with the class embeddings, allowing the model to capture viewpoint transitions at the latent level.
@@
 $\mathbf{f}^{(i)}=\text{Up}_{i}\bigl(\mathbf{f}^{(i-1)}+\text{Proj}_{\mathrm{combined}}\bigl[\boldsymbol{\gamma}(t)\oplus\mathbf{e}_{{C}_{\mathrm{tar}}}\oplus\mathbf{e}_{{c}_{\mathrm{tar}}}\bigr]\bigr).$
-Cross-Attention Module
+3.3.4 Cross-Attention Module
 A cross-attention mechanism is integrated in the mid and up blocks, enabling information flow from the reference to the target view using ray information and latent feature alignment.
 Let $\mathbf{r_{\text{ref}}}$ denote the ray embeddings of the reference view and $\mathbf{r_{\text{tar}}}$ denote the ray embeddings of the target view.
-We use standard ray parameterization as in NeRF (Mildenhall et al.
-2021) to compute ray origins and directions for camera pose encoding to get $\mathbf{r_{\text{ref}}}$ and $\mathbf{r_{\text{tar}}}$.
+We use standard ray parameterization as in NeRF [30] to compute ray origins and directions for camera pose encoding to get $\mathbf{r_{\text{ref}}}$ and $\mathbf{r_{\text{tar}}}$.
 Let $\mathbf{z^{\text{inv}}_{\text{ref},\mu}}$ be the DDIM-inverted latent mean of the reference image, and $\mathbf{f_{\text{tar}}}$ be the intermediate target feature maps at the cross-attention block.
@@
 Using $\tilde{\mathbf{z}}^{\text{inv}}_{\text{tar},\mu}$, we next explain the fusion strategy.
-Fusion Strategy
+3.4 Fusion Strategy
 To synthesize semantically rich target view latents from the predicted DDIM-inverted mean latent $\tilde{\mathbf{z}}^{\text{inv}}_{\text{tar},\mu}$, we introduce two fusion strategies that combine this mean latent with a noise component derived from the input view latent.
 These strategies re‑inject the learned noise variance, that is, the high‑frequency details, into the coarse latent.
-We utilize the fact that the noise/variance term Equation (2) in the DDIM-inverted latent of the input view contains scene-level attributes and characteristics (Staniszewski, Kuciński, and Deja 2025), which can be used to synthesize the scene from a novel view when fused with TUNet’s prediction.
-Strategy A: Variance Fusion via $\sigma$-Component
+We utilize the fact that the noise/variance term Equation (2) in the DDIM-inverted latent of the input view contains scene-level attributes and characteristics [41], which can be used to synthesize the scene from a novel view when fused with TUNet’s prediction.
+3.4.1 Strategy A: Variance Fusion via $\sigma$-Component
 In this strategy, we explicitly extract the variance (or noise) component from the DDIM-inverted latent of the input view, denoted as $\mathbf{z}_{\text{ref},\sigma}^{\text{inv}}$.
@@
 $\mathbf{z}_{\text{noisy}}=\tilde{\mathbf{z}}_{\text{tar},\mu}^{\text{inv}}+\mathbf{z}_{\text{ref},\sigma}^{\text{inv}}.$ (4)
-The fused latent $\mathbf{z}_{\text{noisy}}$ is then passed into the Stable Diffusion U-Net to compute the noise prediction:
-$\boldsymbol{\epsilon}_{\theta}=\text{U-Net}(\mathbf{z}_{\text{noisy}},t).$ (5)
-The initial latent for DDIM sampling is obtained as:
-$\tilde{\mathbf{z}}_{\text{tar}}^{\text{inv}}=\tilde{\mathbf{z}}_{\text{tar},\mu}^{\text{inv}}+\sqrt{1+\alpha_{t+1}}\boldsymbol{\epsilon}_{\theta}.$ (6)
-Strategy B: Direct Noise Addition from Reference Inversion
+The fused latent $\mathbf{z}_{\text{noisy}}$ is then passed into the Stable Diffusion U-Net to compute the noise prediction, $\boldsymbol{\epsilon}_{\theta}=\text{U-Net}(\mathbf{z}_{\text{noisy}},t).$ The initial latent for DDIM sampling is obtained as:
+$\tilde{\mathbf{z}}_{\text{tar}}^{\text{inv}}=\tilde{\mathbf{z}}_{\text{tar},\mu}^{\text{inv}}+\sqrt{1+\overline{\alpha}_{t+1}}\boldsymbol{\epsilon}_{\theta}.$ (5)
+3.4.2 Strategy B: Direct Noise Addition from Reference Inversion
 Here, we directly use the noise component from the full DDIM-inverted latent of the input view $\mathbf{z}_{\text{ref}}^{\text{inv}}$, rather than extracting its variance separately.
 The initial latent $\tilde{\mathbf{z}}_{\text{tar}}^{\text{inv}}$ for DDIM sampling is computed as :
-$\tilde{\mathbf{z}}_{\text{tar}}^{\text{inv}}=\tilde{\mathbf{z}}_{\text{tar},\mu}^{\text{inv}}+\sqrt{1+\alpha_{t+1}}\mathbf{z}_{\text{ref}}^{\text{inv}}.$ (7)
-We generate samples using $\tilde{\mathbf{z}}_{\text{tar}}^{\text{inv}}$ from both Equation (6) and Equation (7), and pick the best generated result where we pick the result which has lower LPIPS score.
+$\tilde{\mathbf{z}}_{\text{tar}}^{\text{inv}}=\tilde{\mathbf{z}}_{\text{tar},\mu}^{\text{inv}}+\sqrt{1+\overline{\alpha}_{t+1}}\mathbf{z}_{\text{ref}}^{\text{inv}}.$ (6)
+We generate samples using $\tilde{\mathbf{z}}_{\text{tar}}^{\text{inv}}$ from both Equation (5) and Equation (6).
 Figure 4: Qualitative results with our 167-class trained model.
-Training Objective
+3.5 Training Objective
 Our training objective is to align the DDIM-inverted latent mean of the prediction and ground-truth.
@@
 $\mathcal{L}_{\text{MSE}}=\left\|\tilde{\mathbf{z}}^{\text{inv}}_{\text{tar},\mu}-\mathbf{z}_{\text{tar},\mu}^{\text{inv}}\right\|_{2}^{2}.$ (7)
-Experiments
-Dataset: MvImgNet (Yu et al.
-2023b) consists of 6.5 million frames of real-world scenes across 238 categories.
+4 Experiments
+Dataset: We perform experiments using MvImgNet [61] and RealEstate10K [66].
+MvImgNet consists of 6.5 million frames of real-world scenes across 238 categories.
 We use two subsets of MVImgNet (i) three scene categories: sofas, chairs, and tables.
-A 90-5-5 (Anciukevičius et al.
-2024) split is used for training, validation, and testing, respectively, determined by lexicographic ordering of the scene identifiers.
+A 90-5-5 [1] split is used for training, validation, and testing, respectively, determined by lexicographic ordering of the scene identifiers.
 (ii) We use 8.5 lakh frames across 167 classes and for each class, we keep 1 scene out of 99 in the test set to evaluate our results and compare with other methods.
+In case of RealEstate10K, we train using 1 million pairs.
+We report additional results of RealEstate10K in the supplementary and demonstrate that our method achieves superior results.
 Pre-Processing: We resize the shorter dimension of the images to be 512 and resize the other dimension to maintain the aspect ratio and then take centre crop of $512\times 512$.
@@
 The dimensions of both class and camera embeddings are 64, and the cross-attention dimension is 768 with an attention head dimension of 64.
-We use the latent diffusion backbone (Rombach et al.
-2022).
+We use the latent diffusion backbone [35].
 For training, we randomly pair frames 1-10 of each scene with frames 15-25.
@@
 We train two models on our subset (i) 3 classes and (ii) 167 classes.
-Our 167 class model is trained for 450 epochs on a single 40GB A400 for 17 GPU days with a batch size of 32 and a learning rate of 1e-5, and we decay the learning rate using a cycle scheduler.
+Our 167 class model is trained for 450 epochs on a single 49 GB RTX A6000 for 17 GPU days with a batch size of 32 and a learning rate of 1e-5, and we decay the learning rate using a cycle scheduler.
 During inference, we generate final results with 30 DDIM sampling steps with the initial latent being Equation (5) or Equation (6), which represents the noisy latent at $t$ = 600.
-We compare our 3-class model with GIBR (Anciukevičius et al.
-2024) and the 167-class model with NViST (Jang and Agapito 2024).
+We compare our 3-class model with GIBR [1] and the 167-class model with NViST [22].
 For GIBR, we have the same train/test split and directly report the results from their paper.
@@
 We report LPIPS and FID scores with the resized results for comparisons.
-We report PSNR and SSIM with the original $512\times 512$ resolution results.
-Quantitative Comparison
+We follow the evaluation protocol as given in [1, 8].
+4.1 Quantitative Comparison
 3-class model: Comparison with GIBR on 3 classes at a resolution of $256\times 256$ is shown in Table 1.
@@
 Similarly, in the case of a bowl (third row, middle column), the shadow is faithfully synthesized.
-Qualitative Comparison
-We compare our results with NViST (Jang and Agapito 2024) in Figure 5.
+4.2 Qualitative Comparison
+We compare our results with NViST [22] in Figure 5.
 It is evident that our method synthesizes the target with high fidelity and is able to generate results for near as well as far target views, where NViST fails.
@@
 Even on unseen classes, we are able to generate high-resolution reconstruction for a diverse set of scenes.
+For evaluation on unseen classes, the unseen class is treated as an additional label.
+We obtain its semantic embedding and provide it to the model at test time.
 Figure 6: Results on 6 unseen classes from MVImgNet
-Out of domain data: To evaluate zero‑shot generalization beyond MVImgNet, we assembled an out‑of‑domain test set by downloading freely‑licensed photographs from Unsplash (Unsplash Contributors 2025) featuring natural scenes.
-We show the results and compare with Zero123++ (Shi et al.
-2023) in Figure 7.
+Out of domain data: To evaluate zero‑shot generalization beyond MVImgNet, we assembled an out‑of‑domain test set by downloading freely‑licensed photographs from Unsplash [47] featuring natural scenes.
+Since web images lack ground-truth camera parameters, we identify the most visually similar scene in MvImgNet in terms of viewpoint.
+The camera parameters of this nearest neighbor are then used as a proxy for the web image.
+For target views, we analogously select the corresponding frame from the same scene in which the closest viewing-angle image resides, and adopt its parameters.
+We show the results and compare with Zero123++ [39] in Figure 7.
 While both methods successfully generate plausible novel viewpoints, our approach produces more faithful surface textures and preserves natural scene characteristics.
 Figure 7: Out-of-domain images
-Ablation
-The ablation results are presented in Table 3.
+4.3 Ablation Study
+Architecture Design: The model design ablation results are presented in Table 3.
 We evaluate the following settings.
@@
 Full model 0.492 15.71 0.523
-Table 3: Ablation study using 3-class.
-Conclusion
+Table 3: Ablation study using 3-class model.
+Fusion Strategy Comparison: We compare perceptual and image quality assessment metrics for the two fusion strategies in Table 4. Variance Fusion achieves better results in all metrics.
+Fusion Method LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$
+Variance (Stgy A) 0.495 15.41 0.521 69.86
+Direct (Stgy B) 0.521 14.76 0.457 102.72
+Table 4: Comparison of perceptual and image-quality assessment metrics between the two fusion strategies, using 3-class model.
+Different Stable Diffusion Pipelines: We evaluate our pipeline using three diffusion backbones.
+Stable Diffusion v1.5 [35] is the pipeline we use by default in all of our experiments.
+Furthermore, we compare our default pipeline with v2.1111https://huggingface.co/stabilityai/stable-diffusion-2-1, and Dreamlike Photoreal 2.0222https://huggingface.co/dreamlike-art/dreamlike-photoreal-2.0.
+As shown in Table 5, performance remains consistent across models, indicating robustness of our framework.
+v2.1 achieves slightly lower FID, reflecting improved generative realism, while v1.5 attains marginally higher PSNR/SSIM.
+Dreamlike Photoreal 2.0 shows increased artifacts, likely due to its photoreal art domain fine-tuning, which makes it perform worse for natural images.
+Pipeline LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$
+Stable Diffusion v1.5 0.491 15.41 0.522 93.57
+Stable Diffusion 2.1 0.491 15.27 0.522 88.95
+Dreamlike Photoreal 2.0 0.507 15.15 0.503 101.84
+Table 5: Comparison across different SD pipelines.
+Different Diffusion Timesteps: We compare the performance at three different timesteps $t=400,600,800$.
+As shown in Table 6, decreasing the diffusion timestep from $t=600$ to $t=400$ leads to a degraded performance as reflected in the scores.
+In our experiments, we observe that at $t=400$, the loss is very high compared to the case of $t=600$.
+This indicates that training is harder for a single-step translation with TUnet for $t=400$.
+At $t=400$, the noise level and the inversion trajectory are insufficient for meaningful variance-based fusion.
+The latent still retains the dominant low-frequency structure, reducing the fusion strategy’s ability to recover and refine high-frequency content.
+The performance is worst at $t=800$ as the inversion at this timestep loses most of its signal to perform any effective translation.
+In contrast, $t=600$ allows a weak yet sufficient signal for effective view translation aided by noise fusion to recover better quality results upon sampling.
+Timestep ($t$) LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$
+400 0.510 15.37 0.528 150.28
+600 0.491 15.41 0.522 93.57
+800 0.550 15.01 0.502 197.13
+Table 6: Comparison at different diffusion timesteps $t$.
+4.4 RealEstate10K
+We compare our results with GenWarp [38] and VIVID [10].
+We use 1K images for testing.
+In Table 1, we can see that our method performs better in terms of LPIPS, PSNR, and SSIM, except for long range LPIPS compared to VIVID.
+Mid-range (30-60 frames) Long-range (60-120 frames)
+Method LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
+VIVID 0.523 13.83 0.439 0.594 12.69 0.410
+Ours 0.503 15.04 0.479 0.609 13.44 0.448
+Table 7: Results on 1K pairs of RealEstate10K. Images are uniformly sampled at random from different scenes.
+5 Conclusion
 In this work, we propose a novel method using TUNet and a fusion strategy to synthesize high-quality novel views.
@@
 Supplementary Material
-Novel View Synthesis using DDIM Inversion
-Supplementary Materials
-Novel View Synthesis and Editing
-As we work with DDIM inverted noise, we have the flexibility to extend our pipeline to editing via textual prompts during the final sampling.
-For our novel view reconstruction, we use BLIP (Li et al.
-2022) generated captions for the input image $\mathbf{x}_{ref}$.
-We edit these prompts to add desired scenarios in our reconstructed novel view.
-Figure 8 shows results of our editing mechanism.
-Figure 8: We use textual prompts to edit the novel view reconstructions as desired.
-In the first row, we place a bag on the bench.
-In the second row, a dog is introduced in the scene.
-In the third row, a pair of shoes are placed by the bench.
-Synthesis of Multiple Views from Single Image
+1 Additional Experimental Results
+We provide additional results on RealEstate10K and MVImgNet, including qualitative ablations and high-resolution evaluations.
+We further study task-level generalization by applying the same DDIM-latent translation and fusion principle to low-light image enhancement, treating it as an image-to-image translation problem.
+1.1 RealEstate10K
+In order to train the network, we sample 1 million source-target pairs.
+We resize the images while preserving the aspect ratio, and then center-crop a region of $\times256$.
+Extrinsics parameters remain the same for all image resolutions, following VIVID.
+To obtain intrinsics, we selected the focal length and principal point based on a resolution of 256 $\times$ 256.
+As RealEstate10K do not have class labels, we do not use class embeddings.
+We follow the evaluation pipeline of VIVID.
+We compute the metrics for $\times256$ resolution.
+To generate the images from GeoGPT, Photometric-NVS, and VIVID, we follow the respective sampling and pre-processing strategy as mentioned in the paper.
+We evaluate on 1K mid-range (30–60 frames) and 1K long-range (60–120 frames) RealEstate10K test pairs.
+All metrics are reported as mean $\pm$ standard deviation over three independently sampled subsets of the official test split.
+We report the results in Table 1.
+While our LPIPS is higher than prior work, we improve PSNR and SSIM over baselines Photometric-NVS and VIVID, especially on long-range pairs where viewpoint extrapolation is most challenging.
+Our method shows (i) best long-range SSIM, (ii) strong PSNR gains over Photometric-NVS and VIVID, and (iii) low variance across sampled subsets.
+Mid-range (30-60 frames) Long-range (60-120 frames)
+Method LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
+GeoGPT 0.318 $\pm$ 0.004 15.863 $\pm$ 0.097 0.474 $\pm$ 0.008 0.409 $\pm$ 0.005 14.025 $\pm$ 0.223 0.416 $\pm$ 0.016
+Photometric-NVS 0.387 $\pm$ 0.004 14.271 $\pm$ 0.023 0.410 $\pm$ 0.002 0.508 $\pm$ 0.001 12.328 $\pm$ 0.086 0.342 $\pm$ 0.009
+VIVID 0.442 $\pm$ 0.070 13.480 $\pm$ 0.309 0.408 $\pm$ 0.027 0.547 $\pm$ 0.040 11.852 $\pm$ 0.733 0.340 $\pm$ 0.060
+Ours 0.510 $\pm$ 0.008 14.902 $\pm$ 0.026 0.471 $\pm$ 0.003 0.609 $\pm$ 0.002 13.243 $\pm$ 0.052 0.430 $\pm$ 0.005
+Table 1: Results on RealEstate10K.
+[…]
+Method #Parameters (M) Inference time ($\times$ Ours)
+[…]
+[…]
+[…]
+[…]
+Table 2: Model size and inference speed comparison on RealEstate10K (relative to Ours). Our method takes 2 seconds.
+In Figure 1, we show the output generated by different methods.
+Compared to prior work, our method produces slightly softer textures but preserves global scene structure and viewpoint alignment.
+This is consistent with our quantitative metrics: despite higher LPIPS, we obtain PSNR/SSIM improvements on challenging long-range pairs, with a compact model that remains efficient at inference.
+Input Target VIVID PhotoNVS GeoGPT Ours
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 1: Qualitative comparison on RealEstate10K.
+1.2 Synthesis of Multiple Views from Single Image
 In Figure 2, we generate multiple frames using a single input image.
@@
 Even for long-range viewpoints, the proposed method achieves good synthesis results.
-Figure 9: Generating multiple frames with single input image.
-Qualitative Ablation Results
+1.3 Qualitative Ablation Results
 In Figure 3, we show qualitative results with our ablation setting.
@@
 Full Model shows the best performance.
-Figure 10: Ablation
-Additional Results
+1.4 Additional Results
 In Table 3, we report the results on the original synthesized resolution of $512\times 512$.
@@
 Table 3: Our results at original resolution of $512\times 512$
+We show additional results using our 3-class and 167-class models in Figures 4 and 5, respectively.
+In Figure 4, our model produces novel views that remain geometrically consistent with the targets while preserving global scene layout.
+Figure 5 further shows improved texture fidelity and sharper high-frequency details.
+1.5 Failure Cases
+In Figure 6, we show the failure cases.
+In the first column, we can see that structure is distorted.
+In the second and third columns, shape is not preserved.
+Similarly, for other columns, we see that texture, count, or shape is not preserved.
+2 Extension to Low Light Image Enhancement Task
+In this section, we show that our method can be extended to image-to-image translation task such as LLIE.
+As LLIE requires high fidelity with respect to input-target pairs, we make use of depth maps.
+The depth maps for input images are obtained using Intel DPT-Large model [^3] https://huggingface.co/Intel/dpt-large.
+As the camera parameters are not available, we do not employ them.
+We evaluate the LLIE application of our proposed method on multiple datasets, including LOLv1 [53], LOLv2 [58], and SICE [4].
+LOLv1 comprises 485 paired low-light and normal-light training images and 15 testing pairs.
+LOLv2 is divided into LOLv2-Real and LOLv2-Synthetic subsets, each containing 689 and 900 training pairs and 100 testing pairs, respectively.
+SICE [4] contains 589 low-light and overexposed images.
+We use 80% of scenes for training and 20% for testing.
+The training image resolution is $\times256$.
+Experiment Settings The diffusion process employs a linear noise schedule with $\beta_{\text{start}}=0.0001$ and $\beta_{\text{end}}=0.02$ over $T=500$ timesteps.
+The model is optimized using AdamW optimizer with learning rate $1\times 10^{-4}$, batch size 16, and trained for 1000 epochs using cosine annealing learning rate scheduling.
+The loss function consists of MSE and LPIPS [64] with weighting factor $\lambda=0.1$ for LPIPS loss.
+Depth conditioning is achieved by concatenating 4-channel low-light latents with single-channel depth maps estimated from enhanced or ground truth images.
+At the test time, we use CIDNet [56] to obtain enhanced images which are used to further obtain depth maps.
+As the number of training images is less, we train using two different settings.
+In the first setting, we combine LOLv1 and LOLv2 dataset.
+As LOLv2 has LOLv2-Synthetic dataset, in order to train only on real datasets, we use a second setting wherein we combine LOLv1, LOLv2-Real, and SICE datasets.
+Evaluation Metrics For quantitative evaluation, we adopt Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [52] as distortion metrics.
+To assess perceptual quality, we report Learned Perceptual Image Patch Similarity (LPIPS) [64] with AlexNet [25] backbone.
+Model evaluation is conducted using DDIM sampling with 50 denoising steps.
+Main Results Table 6 presents the quantitative comparison of our method against state-of-the-art low-light enhancement techniques.
+In case of LOLv2, our method outperforms all the other methods by a huge margin.
+As LOLv2 is larger compared to LOLv1 and a real dataset, the improvement on this dataset is extremely significant.
+In case of LOLv1, our method shows the best LPIPS score.
+To ensure a fair comparison, we also train CIDNet [56] and RetinexFormer [6] using the combined dataset.
+We present these results in Table 6.
+In case of CIDNet, we see that the performance significantly drops in terms of PSNR and LPIPS, though it shows marginal improvement in SSIM.
+The drop in the performance is more pronounced in LOLv2-Synthetic.
+We observe a similar phenomenon in case of RetinexFormer.
+The performance, however, improves in case of LOLv2-Real for both CIDNet and RetinexFormer.
+We further experiment with only real datasets and report results in Table 6.
+We combine LOLV1, LOLV2-Real and SICE.
+We see that for both LOLv2-Real and SICE, our method shows best performance in terms of PSNR.
+In LOLv1, CIDNet performs best.
+We show a visual comparison of our model output with RetinexFormer and CIDNet in Figure 7.
+Even though our method is not explicitly designed for LLIE task, we can see that our model generalizes very well to this task.
+Figure 2: Generating multiple frames with single input image from MVImgNet.
+Figure 3: Qualitative ablation results.
 Figure 4: Qualitative results with our 3-class trained model on MVImgNet test set.
 Figure 5: Additional Qualitative results with our 167-class trained model on MVImgNet test set.
-We show additional results using our 3-class and 167-class models in Figures 11 and 12, respectively.
+Input Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Novel Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Target Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 6: MVImgNet failure cases. Each column represents one scene, while rows correspond to input view, predicted novel view, and target view.
+Methods Color Model Complexity LOLv1 LOLv2-Real LOLv2-Synthetic
+[…]
+SNR-Aware [55] SNR+RGB 4.01 26.35 26.716 0.851 0.152 21.480 0.849 0.163 24.140 0.928 0.056
+Bread [16] YCbCr 2.02 19.85 25.299 0.847 0.155 20.830 0.847 0.174 17.630 0.919 0.091
+PairLIE [13] Retinex 0.33 20.81 23.526 0.755 0.248 19.885 0.778 0.317 19.074 0.794 0.230
+LLFormer [50] RGB 24.55 22.52 25.758 0.823 0.167 20.056 0.792 0.211 24.038 0.909 0.066
+RetinexFormer [6] Retinex 1.53 15.85 27.140 0.850 0.129 22.794 0.840 0.171 25.670 0.930 0.059
+GSAD [20] RGB 17.36 442.02 27.605 0.876 0.092 20.153 0.846 0.113 24.472 0.929 0.051
+CIDNet [56] HVI 1.88 7.57 28.201 0.889 0.079 24.111 0.871 0.108 25.705 0.942 0.045
+Ours RGB 182.04 8.05 27.358 0.848 0.075 28.097 0.810 0.059 19.663 0.671 0.140
+Table 4: LOLv1 and LOLv2 results. Following CIDNet [56], we use GT mean method during testing for LOLv1. Best performance is in bold.
+Methods Color Model LOLv1 LOLv2-Real LOLv2-Synthetic
+[…]
+CIDNet HVI 23.103 0.902 0.106 29.455 0.927 0.071 17.462 0.851 0.201
+RetinexFormer Retinex 22.844 0.827 0.146 28.400 0.877 0.116 16.032 0.749 0.254
+Ours RGB 27.358 0.848 0.075 28.097 0.810 0.059 19.663 0.671 0.140
+Table 5: Results for combined dataset. represents the training on the combined dataset.
+Methods Color Model LOLv1 LOLv2-Real SICE
+PSNR SSIM PSNR SSIM PSNR SSIM
+RUAS [34] Retinex 18.654 0.518 15.326 0.488 8.656 0.494
+LLFlow [51] RGB 24.998 0.871 17.433 0.831 12.737 0.617
+CIDNet [56] HVI 28.201 0.889 24.111 0.871 13.435 0.642
+CIDNet∗∗ HVI 20.560 0.808 22.109 0.841 13.227 0.378
+Ours RGB 26.859 0.845 28.182 0.807 15.554 0.382
+Ours HVI 26.753 0.843 28.072 0.805 15.565 0.390
+Table 6: Results on LOLv1, LOLv2-Real, and SICE datasets. ∗∗ represents the training on the combined dataset.
+Input RetinexFormer RetinexFormer RetinexFormer∗∗ CIDNet CIDNet
+CIDNet∗∗ Ours Ours∗∗ Ours-HVI∗∗ Ground Truth
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Input RetinexFormer RetinexFormer RetinexFormer∗∗ CIDNet CIDNet
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+CIDNet∗∗ Ours Ours∗∗ Ours-HVI∗∗ Ground Truth
+Figure 7: Visual comparison of RetinexFormer, CIDNet, and our method across training configurations.
+Ours, indicates training using LOLv1+LOLv2; ∗∗ indicates training using LOLv1+LOLv2+SICE; Ours-HVI uses HVI color model.